# On the Feasibility of Using LLMs to Execute Multistage Network Attacks

### v2

@@ 6. Implementation
For instance, for the high-level command from an LLM to scan a network, we implement a translation that executes a series of nmap scans. We implement the translations for the lateral movement and escalate privilege actions by manually implementing a small library of mappings of known vulnerabilities and their corresponding exploits. For instance, if an LLM specifies to lateral move into a web server with an Apache Struts vulnerability, Incalmo will identify the vulnerability and execute the low-level exploit.
We also extend Caldera to implement the environment state service and attack graph service. We implement the environment state service’s knowledge base with Python objects. For each low-level Caldera primitive, we create parsers that interpret their output and update the knowledge base. The attack graph service is also implemented in Python and uses the environment state service to determine available attack paths.
@@
Finally, we also extend Caldera to interact with LLMs. We do this by first creating a preprompt that contains: an attacker goal, instructions for formatting queries and actions, a brief documentation about the high-level API, and initial known information about the environment. During the execution phase, we extract the python function between the <action></action> or <query></query> tags. Then Incalmo executes the function and returns the result to the LLM. The LLM will continually execute actions and queries until it specifies a <finished> tag or reaches a time limit.
7. Evaluation
@@
In this section, we conduct end-to-end experiments to show how LLMs can use Incalmo to autonomously conduct multistage attacks. Then, we conduct a factor analysis on the use of Incalmo components (i.e., action planner and services) and LLM model size.
@@
Figure 9. The maximum efficacy across 5 trials. We find that with Incalmo, LLMs can successfully and autonomously conduct multistage attacks in 9 out of 10 environments ranging from 25-50 hosts.
7.1. Setup
First, we use the multistage attack environments from Sec. 2, shown in Table 1. Two of the environments are inspired from public reports of real-world attacks (equifax_report, ; colonial_pipeline_techtarget, ). Two other environments are based on common enterprise network topologies (ciscoEnterpriseNetwork, ; ibmTreeNetwork, ). The remaining environments are inspired by environments in prior work (mirage, ; ringNetwork2013technique, ; ferguson2021_deception_psychology, ; ciscoEnterpriseNetwork, ; dumbbellNetwork, ). Further details about environments can be found in Appendix A.
For these experiments we consider two success metrics:
  • $\bullet$
    Attack success: Our first metric measures the LLM’s ability to achieve their goals. We consider a full success as an LLM achieving all goal states (e.g., exfiltrate all 48 data files in the Equifax-inspired environment) in the attack graph. We consider partial success as the ability to achieve at least one goal state (e.g., exfiltrate at least one data file in the Equifax-inspired environment), and no success as the inability to achieve any goal states.
  • $\bullet$
    Attack graph coverage: Sometimes LLMs are unable to partially succeed, but are still able to make substantial progress in attacking the network. To capture this, We measure the percentage of all states the LLM was able to achieve in the attack graph. These states measure the LLM’s ability at other sub-goals of the multistage attack (e.g., reconnaissance, lateral movement).
@@ 7.1. Setup
We evaluate Incalmo across all 10 environments with six different LLMs of various sizes and from different companies: Sonnet 3.5, Haiku 3.5, GPT4o, GPT4o mini, Gemini 1.5 Pro, and Gemini 1.5 Flash. For each environment, we execute 5 trials for each LLM. We control for the attacker’s budget by setting a time limit for each trial at 75 minutes.
7.2. Results
For the following results, for both types of metrics, we consider the peak or maximum efficacy across trials. We believe this is more realistic as attackers can rerun the LLM on the network multiple times.
Fig. 9 shows the attack success metric of the six LLMs (y-axis) in all environments (x-axis). We order the environments by the most fully successful LLMs, followed by those with the most partially successful LLMs.
@@ 7.2. Results
Finding 1.A: Using Incalmo, LLMs can autonomously and fully succeed at multistage attack in 5 out of 10 environments and partially succeed in 9 out of 10 environments (Fig. 9). In contrast, recall that LLMs without Incalmo, can only partially succeed in 1 out of 10 environments (Fig. 2).
We find that LLMs with Incalmo are able to successfully execute end-to-end multistage attacks across 9 out of 10 of the environments. For instance, in the Equifax-inspired environment, the most realistic environment (equifax_report, ), all six LLMs with Incalmo are able to exfiltrate at least some of the data with four LLMs able to exfiltrate all of the data. In comparison, all LLMs without Incalmo are unable to exfiltrate any of the data.
@@
Figure 10. Maximum percentage of attack graph states achieved by LLMs with and without Incalmo. Across all 10 environments, all LLMs (except Gemini 1.5 Flash) were able to achieve more attack graph states with Incalmo than without Incalmo.
In Fig. 10, we illustrate the maximum percentage of attack graph states achieved by the LLMs in all environments. We compare the number of attack graph states of LLMs with and without Incalmo.
@@
Finding 1.B: Across all 10 environments, LLMs with Incalmo achieve a maximum of 48–100% of attack graph states, whereas without Incalmo, they achieve a maximum of 1–30% (Fig. 10).
@@
LLMs equipped with Incalmo achieved more states in the attack graph, compared to LLMs without Incalmo. LLMs equipped with Incalmo in the worst case environment (Enterprise B with 48% of states achieved) achieved more states than LLMs without Incalmo in the best case environment (Enterprise A with 30% of states achieved). Furthermore, in environments where LLMs without Incalmo achieved a limited number of attack graph states, we can see significant improvements when the LLMs use Incalmo. For instance in the Equifax-inspired environment, 4 different LLMs were able to achieve more than 99% of attack graph states with Incalmo, but without Incalmo, the same 4 LLMs were only able to achieve 0.8% of the attack graph states.
7.3. Factor analysis
Impact of Incalmo modules: First, we assess the relative impact of the two key ideas in Sec. 4: using an action planner to assist LLMs with the implementation of actions and using services to assist LLMs with outputting irrelevant actions.
Action planner Environment and attack graph services
Without Incalmo
Incalmo-WAP
Incalmo-WS
Incalmo
Table 2. We conduct a factor analysis across 4 types of abstractions to determine which abstractions in Incalmo contribute towards enabling LLMs to conduct multistage attacks.
@@ 7.3. Factor analysis
To this end, we conduct a series of ablation experiments, removing different modules as seen in Table 2. First, we create an intermediate version without the action planner, Incalmo-WAP, where LLMs do not have access to the action planner, but can use the environment and attack graph services. Here, LLMs can perform 19 predefined low-level actions, such as reading a file or exploiting Apache Struts. [^9] We require the system to use predefined actions to enable the environment and attack graph services. These low-level actions mirror the library that Incalmo’s action planner uses to translate high-level tasks.
Figure 11. Across all 10 environments, LLMs with Incalmo-WAP performed the worst, no LLMs achieved any part of a goal in all 10 environments. This suggests that Incalmo’s action planner plays a signficant role in enabling multistage attacks.
In Fig. 11, we compare the attack success metric of LLMs equipped with Incalmo-WAP vs.  LLMs equipped with Incalmo. We see a dramatic improvement in attack success when LLMs use Incalmo with the action planner enabled.
@@
Finding 2.A: Across 10 environments, LLMs with Incalmo-WAP were unable to partially complete any goal suggesting that the high-level tasks in Incalmo’s action planner are key for LLMs to execute multistage attacks (Fig. 11).
@@
All 6 LLMs with Incalmo-WAP were unable to partially complete any goal in the 10 environments. In fact, Sonnet 3.5 performed better with command line tools in the 4-Layer chain environment than with Incalmo’s library of low-level actions.
@@
Next, we create a system without the environment and attack graph services, Incalmo-WS, where LLMs still have access to the action planner. Incalmo-WS’s action planner still uses the environment and attack graph services to be environment agnostic, but the services are not accessible to the LLM (unlike Incalmo).
@@
Figure 12. LLMs are able to partially or fully achieve 1–5 more environment goals with Incalmo than with Incalmo-WS. This illustrates that the environment and attack graph services further improves the efficacy of LLMs at conducting multistage attacks.
@@
Now, in Fig. 12, we compare the LLMs equipped with Incalmo-WS to LLMs equipped with Incalmo. Unlike LLMs with Incalmo-WAP, LLMs that use Incalmo-WS are sometimes able to partially and fully succeed at attacks. However, LLMs with Incalmo are able to partially and fully succeed in 1–5 environments.
Finding 2.B: Across all LLMs, LLMs are able to partially or fully achieve 1–5 more environment goals when using Incalmo than with Incalmo-WS illustrating that the environment and attack graph services further improves the efficacy of LLMs at conducting multistage attacks (Fig. 12).
@@
For instance, GPT4o mini with Incalmo-WS only partially succeeds in two environments. In contrast, GPT4o mini with Incalmo is able to partially succeed in all environments and fully succeed in two environments. This suggests that the environment and attack graph services enhance the ability of LLMs to execute attacks
@@
Figure 13. Across 9 out of 10 environments, smaller LLMs with Incalmo partially succeed while larger LLMs without Incalmo only partially succeed in one environment.
@@
Impact of model size: We compare the attack performance of smaller models Incalmo to bigger models without Incalmo. From each LLM company (i.e., Anthropic, OpenAI, and Google), we evaluate their biggest model and smallest model (e.g., GPT4o vs GPT4o mini).
@@
In Fig. 13, we illustrate that smaller LLMs with Incalmo have better attack performance than larger LLMs without Incalmo in 9 out of 10 of the environments. For instance, in the 4-layer chain environment Sonnet 3.5 without Incalmo was able to partially complete a goal by exfiltrating a single file, but Haiku 3.5 with Incalmo was able to exfiltrate all 25 files in the environment. In contrast to conventional wisdom in the LLM world that larger model sizes increase performance (brown2020language, ; kaplan2020scaling, ), we see that in our setting, a good abstraction plays a much bigger role in efficacy.
Reliability of success: Our previous results considered the peak efficacy across trials. A related question is how reliably did attacks succeed? In the interest of brevity, we summarize key results here and show a detailed breakdown in Fig. 14 in Appendix  B. In several environments, several LLMs with Incalmo  reliably achieve partial success. For instance, Haiku 3.5 and Sonnet 3.5 were able to achieve partial success in all 5 trials in 5/10 environments. In regards to full success, LLMs were generally not reliable across all environments. Only Haiku 3.5 in the Colonial Pipeline-inspired environment had full success in all 5 trials and Gemini 1.5 Pro in the Equifax-inspired environment had full success in 4 out of the 5 trials.
8. Discussion and limitations
In this section we discuss some of the key limitations of Incalmo.
Improving partial success: In 4 out of 10 environments, LLMs were only able to partially succeed with Incalmo. Partial success often occurs because LLMs are not persistent in exploring attack paths. For instance, frequently LLMs achieve a single goal and then stop.
However, we note that in many of these trials LLMs could have queried the attack graph service to identify that there were additional paths to explore. We hypothesize that LLMs do not use the full potential of the attack graph service because they have little training data for multistage network attacks and attack graphs. As a result, as future work, we plan to explore adding additional data through fine-tuning LLMs to hopefully improve the ability of LLMs using the attack graph service.
Improving failure scenarios: LLMs with Incalmo were unable to partially succeed in the Enterprise B environment. The Enterprise B environment was the only environment that required both an external scan (to identify vulnerable web servers) and an internal scan (to identify a vulnerable database management server). The best performing LLMs failed to execute the internal scan and were distracted by other potential targets in the network.
We hypothesize that improvements to Incalmo’s attack graph service can help address these types of environments. Incalmo’s attack graph service lack abstraction for fine-grained reasoning about access control between hosts. For instance, a host may only access web services on a server in another subnet because the firewall blocks all other requests. But, hosts on the same subnetwork as the server may be able to access other vulnerable services. Consequently, we believe that extending the attack graph service to reason about fine-grained access control could help LLMs attack environments similar to Enterprise B.

### v3

@@ VI Implementation
For each of the five high-level actions in Sec. IV, we create non-LLM and LLM-based agents that translate the task into low-level primitives (e.g., Python scripts, shell commands). We implement the non-LLM agents for the lateral movement and escalate privilege actions by manually implementing a small library of mappings of known vulnerabilities and their corresponding exploits. For instance, if an LLM specifies to lateral move into a host with a CVE, Incalmo will identify the CVE and execute the low-level exploit.
The LLM-based agents use the Domain-shell prompt and are asked to execute the given high-level action (e.g., exploit a host with a known CVE). The LLM-based agents can either execute bash commands or Python code to accomplish the action. We add the ability to execute Python code because LLMs have shown promise at creating exploits in Python [16].
@@
Finally, we implement code to iteratively prompt LLMs. We do this by first creating a prompt with the onboarding process outlined in Sec. IV. During the execution phase, we extract the python function between the <action></action> or <query></query> tags. Then Incalmo executes the function and returns the result to the LLM. The LLM will continually execute actions and queries until it specifies a <finished> tag or reaches a time limit.
MHBench. We implement MHBench with Python and Ansible code to setup the environments on Openstack. We are working on making MHBench more portable by supporting Docker and other hypervisors.
VII Evaluation
@@
In this section, we conduct end-to-end experiments to show how LLMs can use Incalmo to autonomously conduct multi-host attacks. Then, we conduct a factor analysis of Incalmo components and LLM model size.
@@
Figure 9: The maximum proportion of goal states reached across 5 trials. We find that with Incalmo, LLMs can successfully and autonomously conduct multi-host attacks in 9 out of 10 environments ranging from 25-50 hosts.
@@
Setup. We evaluate Incalmo across all MHBench environments with nine different LLMs of various sizes and from different companies: Sonnet 3.7 Thinking, Sonnet 3.5, Haiku 3.5, GPT4o, GPT4o mini, Gemini 1.5 and 2.5 Pro, and Gemini 1.5 and 2 Flash. For each environment, we execute 5 trials for each LLM. We set a time limit of 75 minutes per trial.
In each trial, we log detailed information about the attack such as raw LLM conversations, attack graph states reached (from the attack graph service), tasks executed, and the events from tasks. We use these logs to calculate MHBench metrics. For instance, to calculate percentage of attack graph states reached we compare the attack graph states reached to the environment’s reference attack graph.
For attack success, we report the peak or maximum success across trials; this simulates a realistic model where attackers can rerun the LLM-Incalmo workflow until they succeed once.
VII-A Attack success evaluation
@@
Finding 1.A: Using Incalmo, LLMs can autonomously and fully succeed at multi-host attack in 5 out of 10 environments and partially succeed in 9 out of 10 environments (Fig. 9). In contrast, LLMs without Incalmo do not fully succeed in any environment.
LLMs with Incalmo are able to partially succeed in 9 out of 10 of and completely succeed in 5 out of 10 environments as seen in Fig. 9. We order the environments by the most fully successful LLMs, followed by those with the most partially successful LLMs. For instance, in the Equifax-inspired environment [37], all six LLMs with Incalmo are able to exfiltrate some of the data, with four able to exfiltrate all of the data.
@@
Figure 10: Maximum percentage of attack graph states reached by LLMs with and without Incalmo. Across all environments, LLMs were able to reach more relevant states with Incalmo than without Incalmo.
@@
Finding 1.B: Across all 10 environments, LLMs with Incalmo reach a maximum of 48–100% of attack path states, whereas without Incalmo, they reach a maximum of 1–30% (Fig. 10).
@@
Fig. 10 shows that LLMs using Incalmo reached more states on attack paths, compared to LLMs without Incalmo. LLMs equipped with Incalmo in the worst environment (Enterprise B with 48% of states reached) reached more states than LLMs without Incalmo in the best case environment (Enterprise A with 30% of states reached, see Fig. 4). Furthermore, in environments where LLMs without Incalmo reached few relevant states, we can see significant improvements when the LLMs use Incalmo. For instance in the Equifax-inspired environment, four LLMs were able to reach more than 99% of attack path states with Incalmo, but without Incalmo, the same four LLMs were only able to reach 0.8% of the relevant states.
We only evaluate attack graph states reached for six of the nine LLMs because of the manual effort required for determining attack graph states reached without Incalmo (see Sec. III). In particular, we add small LLMs for each provider as a reference for a future evaluation in Sec. VII-C.
Robustness of success. Our previous results considered the peak efficacy across trials. A related question is how reliably did attacks succeed? In the interest of brevity, we summarize key results here and show a detailed breakdown in Fig. 15 (in Appendix  B). In several environments, several LLMs with Incalmo  reliably achieve partial success. For instance, in the Equifax environment Haiku 3.5 and Sonnet 3.5 were able to achieve partial success in all 5 trials. With regards to full success, the story is more mixed. Only Haiku 3.5 in the Colonial Pipeline-inspired environment had full success in all 5 trials and Gemini 1.5 Pro in the Equifax-inspired environment had full success in 4 out of the 5 trials.
Figure 11: Minutes taken for fully successful attacks. Attacks range from taking 14 to 70 minutes.
VII-B Cost and time metrics
Next, we measure the time and cost of LLMs using Incalmo to execute attacks. In terms of time, LLMs with Incalmo can rapidly attack networks. For instance, in the Colonial Pipeline-inspired environment, LLMs are able to use Incalmo to successfully infect all 15 target hosts in 12 to 18 minutes (Fig. 11). Similarly, LLMs can use Incalmo to exfiltrate all 48 databases in the Equifax-inspired environment in just 54 minutes.
However, sometimes LLMs were inefficient in there attacks. For instance, in one trial of Dumbbell A, Haiku 3.5 took 35 extra minutes because it infected all 15 external web servers twice. And only after, Haiku 3.5 with Incalmo infected and exfiltrated data from database instances.
TABLE IV: Token Cost of Multi-Host Attacks in 1,000s of Tokens
Environment Input Tokens Output Tokens
Min Mean Max Min Mean Max
GPT4o mini $4.1$ $104.4$ $1474.3$ $0.2$ $0.2$ $15.6$
GPT4o $9.4$ $106.5$ $1005.7$ $0.9$ $3.2$ $11.9$
Gemini 1.5 Flash $3.5$ $9.8$ $26.1$ $0.2$ $0.2$ $0.7$
Gemini 2 Flash $12.2$ $137.2$ $1189.1$ $1.0$ $3.0$ $10.9$
Gemini 1.5 Pro $6.7$ $29.1$ $243.2$ $0.3$ $1.0$ $4.4$
Gemini 2.5 Pro $7.4$ $672.4$ $2022$ $0.9$ $9.7$ $19.8$
Haiku 3.5 $14.6$ $799.2$ $4241.7$ $1.4$ $12.5$ $50.9$
Sonnet 3.5 $57.5$ $862.8$ $5897.1$ $5.0$ $19.3$ $60.1$
Sonnet 3.7 $61.0$ $279.3$ $997.8$ $2.5$ $6.0$ $19.6$
In addition, these attacks are relatively inexpensive. For instance, the Gemini 2 Flash usage is within Google’s free usage tier using only a max of 1,220K input tokens and 0.7K output tokens, while the free tier supports up to 250K tokens per minute. In Table IV, we show how the most expensive attack was Sonnet 3.5 with a max of 5,750K input tokens and 60K output tokens; or around \$15.
Taken together, the attack success, the low cost, and quick time results present a significant milestone in using LLMs for realistic cyber-offense capabilities. For defenders, enterprises conducting penetration tests often invest significant resources to hire domain experts who attack and uncover hidden threats. These results highlight the potential for LLMs to significantly reduce the cost of these penetration tests. On the attack side, at the same time, it also raises the threat of malicious adversaries launching low-cost complex attacks against enterprise targets.
VII-C Factor analysis
Next, we conduct a series of ablation experiments, varying several factors: (1) disabling different modules in Incalmo; (2) varying the size of the models; and (3) using alternative LLM-based realizations of the Incalmo high-level actions. In the interest of brevity, we focus on a smaller set of LLMs that were successful in the earlier end-to-end results.
@@ VII-C Factor analysis
Impact of high-level actions. First, we create an intermediate version without high-level actions, Incalmo-WHT, where LLMs do not have access to the five high-level actions, but can use the environment and attack graph services. Here, LLMs can perform 19 predefined low-level actions, such as reading a file or exploiting Apache Struts. [^11] We require the system to use predefined actions to enable the environment and attack graph services. These low-level actions mirror the library that Incalmo’s uses to translate high-level tasks.
@@
Finding 2.A: Across 10 environments, all LLMs were unable to partially complete any goal without Incalmo’s high-level actions, suggesting that the abstraction is an important factor for attack success.
@@
All nine LLMs with Incalmo-WHT were unable to partially complete any goal in the 10 environments. In fact, Sonnet 3.7 Thinking performed better with the command line tools in Sec. II because it had partial success in the 4-Layer chain environment unlike Sonnet 3.7 Thinking with Incalmo-WHT.
@@
Impact of Incalmo services. Next, we create a variant of Incalmo without the environment and attack graph services called Incalmo-WS, but LLMs still have access to the five high-level actions. Incalmo-WS’s high-level agents still use the environment and attack graph services to be environment agnostic, but the services are not accessible to the LLM (unlike Incalmo).
@@
Figure 12: LLMs are able to partially or fully achieve 1–5 more environment goals with Incalmo than with Incalmo-WS. This illustrates that the environment and attack graph services further improves the efficacy of LLMs at conducting multi-host attacks.
@@
Now, in Fig. 12, we compare the LLMs equipped with Incalmo-WS to LLMs with Incalmo.121212 For this analysis, we were unable to show the new Gemini 2 models because the models are recently released and have strict usage limits. We plan to run the experiments and add the data. Unlike LLMs with Incalmo-WHT, LLMs that use Incalmo-WS are sometimes able to partially and fully succeed at attacks. However, LLMs with Incalmo are able to partially and fully succeed in 1–5 environments.
Finding 2.B: LLMs are able to partially or fully succeed in 1 to 5 more environments when using Incalmo vs.  disabling the environment and attack graph services. This suggests these Incalmo services are also important for attack success (Fig. 12).
@@
For instance, GPT4o mini with Incalmo-WS only partially succeeds in two environments. In contrast, GPT4o mini with Incalmo is able to partially succeed in all environments and fully succeed in two environments. This suggests that the environment and attack graph services enhance the ability of LLMs to execute attacks.
@@
Figure 13: Across 9 out of 10 environments, smaller LLMs with Incalmo partially succeed while larger LLMs without Incalmo only partially succeed in one.
@@
Impact of model size. We compare the attack success of smaller models with Incalmo to bigger models without Incalmo. From each LLM vendor (e.g., OpenAI), we evaluate a small and big model (e.g., GPT4o vs GPT4o mini).
@@
In Fig. 13, we illustrate that smaller LLMs with Incalmo have better attack performance than larger LLMs without Incalmo in 9 out of 10 of the environments. For instance, in the 4-layer chain environment Sonnet 3.5 without Incalmo was able to partially complete a goal by exfiltrating a single file, but Haiku 3.5 with Incalmo was able to exfiltrate all 25 files in the environment. In contrast to prevailing wisdom that larger model sizes are more critical for performance [10, 28], in our cyber offense domain, we see the Incalmo abstraction and services are more critical than model size.
Figure 14: The maximum proportion of goals Sonnet 3.5 agents were able to achieve in three environments. Sonnet 3.5 agents show promise at individual tasks, but LLMs still require assistance from Incalmo agents to successfully execute attacks. The gray boxes are environments where that task is not required to achieve the goals.
Using LLM-based agent realizations of high-level actions. In the prior evaluations, high-level actions are executed by our stateful non-LLM agents. Next, we substitute Incalmo’s action agents (e.g., Incalmo’s scan agent) with LLM-based agents. For example, when an LLM planner initiates a lateral movement action, instead of Incalmo’s agent retrieving a predefined exploit, a separate LLM-based agent dynamically generates the exploit by writing Python code.
For this analysis, we evaluate Sonnet 3.5 as the “planner”’ agent and for “execution” agents in three environments: Equifax-inspired, Colonial Pipeline-inspired, and 6-layer star. We select Sonnet 3.5 as it was the leading LLM that performed well with and without Incalmo. Each LLM-based execution agent had a limit of 10 interactions for each high-level action131313 We limit the interactions and only evaluate the LLM-agents in three environments because of the cost per trial, each trial with agents can cost up to \$30, while the trials without agents were \$0.1 - \$15..
We consider two experimental setups: (1) all execution agents use Sonnet 3.5 instead of Incalmo and (2) replace Incalmo execution agents one at a time.
Finding 4: Sonnet 3.5-based action execution agents show promise at executing lateral movement, network scanning, privilege escalation, and data exfiltration. But LLMs still require assistance from Incalmo agents to successfully execute attacks (Fig. 14).
In the “all” setup, Sonnet 3.5 as the planner only using Sonnet 3.5 execution agents fails to achieve even partial success across the three evaluated environments, as seen in Fig. 14. However, when replacing a single Incalmo agent for LLM-based agents, Sonnet 3.5 can partially succeed in all three environments depending on the specific type of agent. For instance, Sonnet 3.5 with a Sonnet 3.5 lateral movement agent (with the others being Incalmo agents) was able to partially succeed in all three environments.
This study serves three additional purposes. First, it identifies specific high-level steps LLMs may have struggled earlier without Incalmo. (We could not have done this earlier since the LLM failed early and never got to later stages in the attack graph.) Second, it further demonstrates the modularity and extensibility of Incalmo to incorporate new execution agents. Third, it also suggests a viable roadmap for Incalmo to tackle unforeseen scenarios via AI-based action agents as a fallback when its built-in agents may lack coverage.
VIII Discussion and limitations
Improve partial success. In 4 out of 10 environments, LLMs were only able to partially succeed even with Incalmo. This occurs because LLMs often achieve a single goal and then stop. We noticed that in many of these trials LLMs could have queried the Incalmo attack graph service to identify that there were additional paths to explore. We speculate this could be because LLMs may not have much training data for multi-host attacks and using attack graphs. An interesting direction for future work is to further train and fine-tune LLMs to better leverage the attack graph service.
Improving failure scenarios. In Enterprise B, all LLMs failed using Incalmo. It turns out, this setting required both an external scan (to identify vulnerable web servers) and an internal scan (to identify a database management server). Even the best performing LLMs failed to execute the internal scan and were distracted by other potential targets. We hypothesize that improvements to Incalmo’s attack graph service can help address these types of environments. Incalmo’s current attack graph service implementation does not have features to reason about network segments and access control (e.g., need to be on same subnet to access a host because of inter-segment firewalls). We believe that extending the attack graph service to reason about fine-grained access control could further improve attack success for settings similar to Enterprise B.
Extending Incalmo capabilities. For instance, if an environment has a vulnerability but the Incalmo lateral movement agent does not have an exploit, then it will currently fail. That said, Incalmo is extensible and we can consider two ways to extend Incalmo: 1) community developers add the exploit to the Incalmo knowledge base, or 2) use the LLM-based agents as seen in Sec. 14 to try to dynamically generate the exploit.